OpenStudy (anonymous):

Find the function value. f(x)=x^3 f(-9)=___

OpenStudy (anonymous):

ok so its asking you this: f(when x is 9) = \[x^{3}\]

OpenStudy (anonymous):

what does x^3 equal when x=9

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

or -9

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

so whats the problem?

OpenStudy (anonymous):

-9, sorry

OpenStudy (anonymous):

so can you compute this?

OpenStudy (anonymous):

i am not sure, first time taking algebra 1

myininaya (myininaya):

\[f(a)=a^3\] \[f(5)=5^3\] so f(-9)=?

OpenStudy (anonymous):

Myininaya..now i am confused!

OpenStudy (anonymous):

where does the 5 come from?

myininaya (myininaya):

its an example

OpenStudy (anonymous):

oh, ok..

myininaya (myininaya):

\[f(6)=6^3\] \[f(-6)=(-6)^3\]

OpenStudy (anonymous):

ok something cubed looks like \[ a^{3} = a\times a \times a\]

myininaya (myininaya):

f is like a maching you are putting a number into that machine and you are getting the input cubed back

myininaya (myininaya):

machine*

OpenStudy (anonymous):

so \[x^{3} = x \times x \times x \]

OpenStudy (anonymous):

now let each x = -9 and multiply that together.

OpenStudy (anonymous):

so i would multiply-9 *3?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

-27

OpenStudy (anonymous):

wait, I never said there was a three involved, did I?

OpenStudy (anonymous):

no

myininaya (myininaya):

\[(-9)^3\]

myininaya (myininaya):

\[(-9)^3 \neq -9 \cdot 3\]

OpenStudy (anonymous):

if you were to let each of those x's that i wrote out earlier what would that look like when you substitute them in?

OpenStudy (anonymous):

let each of those = 9 i meant.

myininaya (myininaya):

-9*

OpenStudy (anonymous):

just like myininaya did, thats what i meant..

myininaya (myininaya):

\[\text{ what is } (-9)^3 \text{?}\]

OpenStudy (anonymous):

9*3=27

myininaya (myininaya):

no

OpenStudy (anonymous):

yes -9 is correct so let me start it for you f(-9) = \[-9 \times -9 \times ? \]

OpenStudy (anonymous):

what else goes there to finish what i wrote out before?

myininaya (myininaya):

example: \[a^3=a \cdot a \cdot a\] \[5^3=5 \cdot 5 \cdot 5 =25 \cdot 5 = 125\]

OpenStudy (anonymous):

-9x-9x-9=729

myininaya (myininaya):

almost

OpenStudy (anonymous):

wait what is -1 x -1 ?

myininaya (myininaya):

a negative times a negative times a negative =?

OpenStudy (anonymous):

-1

myininaya (myininaya):

\[-1 \cdot -1=1\]

myininaya (myininaya):

\[-1 \cdot -1 \cdot -1=1 \cdot -1 =?\]

OpenStudy (anonymous):

ugh, i get mixed up with these negatives..sorry for being difficult

OpenStudy (anonymous):

-1 * -1 = ? then whatever ? is needs to be multiplied by -1 so ? * -1 = your sign answer

OpenStudy (anonymous):

so f(-9)=9

OpenStudy (anonymous):

sorry, still confused, and this is probably a simple problem and I am just not getting it

OpenStudy (anonymous):

wait what happened? i thought you got the correct solution above but just the wrong sign?

OpenStudy (anonymous):

am i suppost to put a negative before the 729.....-729?

OpenStudy (anonymous):

you need to re-read what we talked about above and decide for yourself...ask any questions you have about clarity but just re-read what is above the answer is there!

OpenStudy (anonymous):

wasn't getting it but thanks.

OpenStudy (anonymous):

ok, so moses, this i show i understood it -9*-9=81*-9=-729

OpenStudy (anonymous):

you are intelligent!!!!

OpenStudy (anonymous):

good job that is right!

OpenStudy (anonymous):

thanks again..i had to read it twice. I think what confused me more is when the other person came on here and it really through me off..so sorry for the confusion

OpenStudy (anonymous):

it is -729